OpenStudy (ashy98):

12 subtracted from x is equal to x divided by 4. What is x???

OpenStudy (anonymous):

Ok now we need to try to simplify this so I would multiply both sides by 4

OpenStudy (anonymous):

do that and tell me your result

OpenStudy (ashy98):

8?

OpenStudy (anonymous):

Well you distribute 4(12-x)=x and that makes it 96-4x=x

OpenStudy (anonymous):

since you multiply 4 on both sides but now just do add 4x on both sides. So you cancel the 4x on the left side then add 4x to the right side and you get 96=5x

OpenStudy (anonymous):

Now divide by 5 on both sides \[\frac{ 96 }{ 5 } = \frac{ 5x }{ 5 }\]

OpenStudy (anonymous):

Tell me what you get

OpenStudy (anonymous):

x=16